OpenStudy (anonymous):

What are the x and y intercepts for the graph of the equation 5x – y = 20?

OpenStudy (raden):

hint : to get x-intercept take y=0, solve for x to get y-intercept always take x=0, solve for y

OpenStudy (anonymous):

so x=4 right?

OpenStudy (raden):

yes, completely (4,0)

OpenStudy (raden):

next, y-intercept ?

OpenStudy (anonymous):

y=20?

OpenStudy (raden):

5x-y=20 (plug x=0) 5(0)-y=20 0-y=20 -y=20 multiply by (-1) to both sides (-1) * -y = (-1) * 20 y = -20

OpenStudy (anonymous):

O ok

OpenStudy (raden):

completely (0,-20) for y-intercept

OpenStudy (anonymous):

So my answer is (4,-20) right?

OpenStudy (raden):

no,, x-intercept different with y-intercept .... we got x-intercept is (4,0) and y-intercept is (0,-20)

OpenStudy (anonymous):

O ok.... Thank you!!!

OpenStudy (raden):

welcome